OpenStudy (anonymous):

Find the center (h,k) and the radius of a circle 5x^2-8x+ 5y^2-5y-5= 0 ??

OpenStudy (anonymous):

start with \[5x^2-8x+5y^2-5y=5\] then divide by 5 to get \[x^2-\frac{8}{5}x+y^2-y=1\] the "complete the square"

OpenStudy (anonymous):

\[(x-\frac{4}{5})^2+(y-\frac{1}{2})^2=1+\frac{16}{25}+\frac{1}{4}\] \[(x-\frac{4}{5})^2+(y-\frac{1}{2})^2=\frac{189}{100}\]

OpenStudy (anonymous):

center is \[(\frac{4}{5},\frac{1}{2})\] radius is \[\frac{\sqrt{189}}{10}\]

OpenStudy (anonymous):

Thanks a lot!!

OpenStudy (anonymous):

yw